• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Moffit

    Log entry time 19:46:40 on June14,2004

    Entry number 125756

    keyword=Septum readback frozen... No detector signals

    Noticed that there was no change in the detector signal with beam on... I was just seeing pedestals. Determined that the septum controls have not been reading back for the last 30 minutes... and could be off.

    Paging Tech On call.